I can't for the life of me solve this equation. Could you please show me the steps?
m+6sqrt(m)-16=0

Answer by stanbon(75887)

m+6sqrt(m)-16=0
----
It doesn't look like it, but this is a quadratic
equation where sqrt(m) is the variable.
------
(sqrt(m))^2 + 6sqrt(m) - 16 = 0
-----------------------------------
Factor:
(sqrt(m)+8)(sqrt(m)-2) = 0
-------
sqrt(m) = -8 or sqrt(m) = 2
-----
Real Number solution:
m = 2^2 = 4
--------------------
Cheers,
Stan H.
===================